OpenStudy (anonymous):

Write the equations of two tangent lines y=4/x through P(3,1) (P is not on f(x) graph

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

well its easy let solve it . let find the slope of tangents . how do you find the slope of tangent to a curve y=4/x?

OpenStudy (anonymous):

I know I have to find the derivative to find the slop of the function

OpenStudy (anonymous):

but that's kind of it haha

OpenStudy (anonymous):

okay lets find the slope first .. for at any point x on curve y= 4/x slope of tangent at that point will be dy/dx so take derivative with respect x of given equation can you find that?

OpenStudy (anonymous):

-4x^-2

OpenStudy (anonymous):

right so now slope of tangent at point x will -4/x^2 and we have a point (3,1) so now can you find equation of tangent by slope point form?

OpenStudy (anonymous):

is it y-1= -4/x^-2 (x-3) ? I'm not sure

OpenStudy (anonymous):

if I plug x values into the original equation and then put these x and y values into the tangent line, I just get something like 5=5

OpenStudy (anonymous):

i don't really get how to find the equations of the 2 tangent lines for the graph from here

OpenStudy (anonymous):

okay just wait for a sec

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

okay i made a big mistake sorry for that really.. we have to find a tangent to given a curve which will pass through the point (3,1) so we find slope of tangent which is equal to -4/x^2 so we have equation of tangent as y-1= -4/x^2 (x-3) we will write this as y-1=m(x-3) where m=-4/x^2 now as we want to find a point which on the graph and through we draw tangent which will pass through point (3,1) so we will put y=4/x in equation (1) (remember why we did that because its an important step) so we get (4/x)-1=-4/x^2 (x-3) by simplifying this we get a quadratic equation x^2-8x+12=0 by solving that we get x=2 and x=6 now put this values of x in slope of tangent we get slope as -1 and -1/9 so now putting values of slope in equation of tangent we get for m=-1 y-1=-1(x-3) y-1=-x+3 x+y-4=0 and for m=-1/9 we have y-1=-1/9(x-3) we get 9y-9=-x+3 that is x+9y-12=0 this is your answer,,!! :)

OpenStudy (anonymous):

okay i get it now :) thank you so much this is so helpful!

OpenStudy (anonymous):

you are most welcome ..! :)